OpenStudy (anonymous):

How many solutions (using only nonnegative integers) are there to the following equation? x1 + x2 + x3 + x4 + x5 = 31

OpenStudy (anonymous):

I need an numerical number. Infinite doesn't work

OpenStudy (loser66):

what do x1 ,x2 ,....mean? \(x^2\)?

OpenStudy (loser66):

or \(x_2\)? just subscript?

OpenStudy (anonymous):

I guess it's written as x

OpenStudy (anonymous):

From what i understand its essentially 31+0+0+0+0=31 30+1+0+0+0=31 29+1+1+0+0=31 so on and so forth

ILovePuppiesLol (ilovepuppieslol):

we need 5 numbers which add up to 31 we can also say that we need to divide 31 into 5 parts these are 31 blocks- \(\square \square\square\square\square\square\square\square\square\square\square\square\square\square\square\square\square\square\square\square\square\square\square\square\square\square\square\square\square\square\square\) we need to find number of ways in which we can divide these blocks into 5 parts. one such way is this- \(\square \square\square {\Huge\color{red}|}\square\square\square\square\square{\Huge\color{red}|}\square\square\square\square\square{\Huge\color{red}|}\square\square\square\square\square\square\square\square\square\square{\Huge\color{red}|}\square\square\square\square\square\square\square\square\) the 4 red sticks are dividing the 31 blocks into 5 parts :) the drawing above is just a case i took, we can change the position of sticks and they will always divide the 31 blocks into 5 parts so we can say that the total number of solutions is equal to the total number of ways in which we can distribute the sticks in this system- \(\square \square\square {\Huge\color{red}|}\square\square\square\square\square{\Huge\color{red}|}\square\square\square\square\square{\Huge\color{red}|}\square\square\square\square\square\square\square\square\square\square{\Huge\color{red}|}\square\square\square\square\square\square\square\square\) there are 31 blocks so therefore there are 30 spaces in between them where we can place the 4 sticks so total number of ways must be this-> \(\Large ~^{30}C_4\)

OpenStudy (anonymous):

\[\left(\begin{matrix}30 \\ 4\end{matrix}\right) = ? \] This is where I got but am lost past this

OpenStudy (anonymous):

I think thats what you meant by 30 C 4

ILovePuppiesLol (ilovepuppieslol):

:D yes thats what i meant \(\left(\begin{matrix}30 \\ 4\end{matrix}\right) \) can be written like this- \(\left(\begin{matrix}30 \\ 4\end{matrix}\right) = \large \frac{\lfloor{30} \rfloor}{\lfloor{4} \rfloor \lfloor{30-4} \rfloor}\)

OpenStudy (anonymous):

How do I come to a solution from there?

ILovePuppiesLol (ilovepuppieslol):

do you know about factorials?

ILovePuppiesLol (ilovepuppieslol):

\(\lfloor{n}\rfloor\) denotes \(\color{blue}{n ~factorial}\) we can also write it like this- \(\lfloor{n} \rfloor=n!\) ^this is just a different notation of factorial thingy what "n factorial" is really-> \(n!=n \times (n-1) \times (n-2) \times ...... (3) \times (2) \times (1)\) its just the multiplication of all numbers starting from 1 till n :)

ILovePuppiesLol (ilovepuppieslol):

all *natural numbers

OpenStudy (anonymous):

So to solve \[30i \div 4i(30i-4i) \] or \[30i \div (4i \times 26i)\]

ILovePuppiesLol (ilovepuppieslol):

its not "\(i\)" its denoted by the exclamation mark "\(!\)" and it can be solved like this- \(\Large \frac{30!}{4!(30-4)!}\) =\(\Large \frac{30!}{4!(26)!}\) =\(\Large \frac{30 \times 29 \times ..... \times \color{red}{ 26 \times 25 \times......2 \times 1}}{4! {(\color{red}{26 \times 25 \times...... \times 2 \times 1})}}\) the red part from numerator and denominator gets cancelled and you get this- =\(\Large \frac{30 \times 29 \times 28 \times 27}{ 4 \times 3 \times 2 \times 1}\) now from here you must simplify this and get the answer =]

OpenStudy (anonymous):

\[657720\div24=27405\] When I put that in it shows incorrect

ILovePuppiesLol (ilovepuppieslol):

this calculation seems length tho xP you can use a calculator if its allowed \(\Huge \ddot \smile\)

OpenStudy (anonymous):

sorry I dont know I kept putting i instead of !

OpenStudy (anonymous):

Am I doing the math incorrectly or missing a step?

ILovePuppiesLol (ilovepuppieslol):

no you did it correct i made a mistake in this part- http://prntscr.com/ahty2q it must be this- there are 31 blocks and 4 sticks so total number of ways to distribute 31 into 5 parts must be equal to total number of combinations of these 31 blocks and 4 sticks so total number of combinations of these 31 blocks and 4 sticks = \(\large ~^{35}C_4\)

OpenStudy (anonymous):

So does that just change the equation to 35!/4!(30-4)! or (35-4)!

ILovePuppiesLol (ilovepuppieslol):

it changes it to this- \(\large \frac{35!}{4!(35-4)!}\)

OpenStudy (anonymous):

Came to 52360. That correct?

OpenStudy (anonymous):

35*34*33*32=1256640 1256640/24=52360

ILovePuppiesLol (ilovepuppieslol):

yeah that looks alright to me

OpenStudy (anonymous):

Perfect thank you very much

ILovePuppiesLol (ilovepuppieslol):

:'D no prblem lol